What is the next term of the sequence 9, 6, 4, 8/3, . . . ?

Found 2 solutions by Pieter, junit_2010:
Answer by Pieter(35)

if you look at the sequence:
to get from 9 to 6: you have to multiply by 2/3.
Now you aren't sure if that is the mean operation you have to do.
To be sure: just check on the other numbers if you are at the right end:
And yes: 6*(2/3)=4 ; 4*(2/3)=8/3.
So the next term in the sequence will be (8/3)*(2/3)=16/9.
This is an example of a GEOMETRIC SEQUENCE.
Greets, Pieter

Answer by junit_2010(2)

6*(2/3)=4 ; 4*(2/3)=8/3.
So the next term in the sequence will be (8/3)*(2/3)=16/9.